OpenStudy (anonymous):

Can someone help with the set up of this? Find the flux of the vector field F=(5x-y)i-2yj+7yzk out of the region W which is defined by the surfaces x=y^2, x=4, z=0, and x=z

OpenStudy (anonymous):

Well, to calculate flux, we can do the integral: \[\int\limits del . F(x,y,z) dA\] (sorry, there doesn't seem to be a del symbol available, and that's a dot product between del and F. You may know it as "divergence") Do you know how to use this?

OpenStudy (anonymous):

My main problem is trying to set up the parameterization, i don't know if there is a way to do it with just one. Also, looking at the surface, it is akward to try and figure out the bounds for u and v.

OpenStudy (anonymous):

I made a mistake, anyway. Integral should be over dV :)

OpenStudy (anonymous):

We can use the Divergence Theorem to change the integral to one over an area. Can you show me exactly where you're at in the whole process?

OpenStudy (anonymous):

I have the integral for the first surface, the flat vertical one, \[\int\limits_{0}^{4}\int\limits_{-2}^{2}(20-u) du dv = 320\] Which I think is right....but not totally sure.

OpenStudy (anonymous):

Hmm, trying to see exactly what you did.

OpenStudy (anonymous):

I had the parameterization \[\Phi(u,v)=<4,u,v> \] with u from -2 to 2 and v from 0 to 4. then I took the div of F to be 5-2+7y and substituted in u for y, then took the dot product of that and the cross product of \[\Phi _{u} and \Phi _{v}\]

OpenStudy (anonymous):

Ah, I see. I'll be honest, I'm rusty with approaches like this. I'd probably advise trying to get a hold of someone who is less rusty. : /

OpenStudy (anonymous):

Thanks, I figured out that I was doing it wrong anyways.

OpenStudy (anonymous):

Trying to get someone else for you, all the same :)